# Learning topological states from randomized measurements using variational tensor network tomography

### v3

@@ Abstract
Learning faithful representations of quantum states is crucial to fully characterizing the variety of many-body states created on quantum processors. While various tomographic methods such as classical shadow and MPS tomography have shown promise in characterizing a wide class of quantum states, they face unique limitations in detecting topologically ordered two-dimensional states. To address this problem, we implement and study a heuristic tomographic method that combines variational optimization on tensor networks with randomized measurement techniques. Using this approach, we demonstrate its ability to learn the ground state of the surface code Hamiltonian as well as an experimentally realizable quantum spin liquid state. In particular, we perform numerical experiments using MPS ansätze and systematically investigate the sample complexity required to achieve high fidelities for systems of sizes up to $48$ qubits. In addition, we provide theoretical insights into the scaling of our learning algorithm by analyzing the statistical properties of maximum likelihood estimation. Notably, our method is sample-efficient and experimentally friendly, only requiring snapshots of the quantum state measured randomly in the $X$ or $Z$ bases. Using this subset of measurements, our approach can effectively learn any real pure states represented by tensor networks, and we rigorously prove that random-$XZ$ measurements are tomographically complete for such states.
@@ I Introduction
Tomographic and other recovery methods for extracting information from these experiments are crucial, in particular, when they provide detailed diagnostic information about the prepared quantum state [11]. Indeed, recent approaches using randomized measurements have shown promise for measuring specific properties, such as the Renyi entropy [8, 12, 13, 14]. However, to extract Von Neumann entanglement entropy across arbitrary subsystem sizes using these methods might require multiple copies, substantially more experimental control, or more samples than what is currently available. Quantum state tomography (QST) is designed to reconstruct a quantum state from informationally complete measurements and is useful for systems where the desired observables are not known a priori [15] or cannot be measured due to the constraints of the experimental setup. Although QST is generally not scalable [16, 17], it becomes significantly more effective when applied with priors for particular classes of states. This makes QST a valuable alternative for extracting many properties from such systems, particularly when traditional methods are limited. Established state tomographic methods like matrix product state (MPS) tomography [18, 19, 20] provide a natural description with robust guarantees for one-dimensional systems but fall short in accurately capturing two-dimensional quantum states, especially when they exhibit non-injective properties [21] and complex entanglement patterns [22]. While neural network quantum state tomography (NNQST) [23, 24] offers yet another option, its efficacy for topological states also remains unclear due to their intrinsic long-range entanglement and intricate phase structures across different components of the wavefunction [25].
@@
Figure 1: Schematic illustration of variational tensor network tomography. Data collection: A target state vector $|\phi\rangle$ is measured in multiple bases for a collection of $N$ samples defined in  Eq. (1). For instance, the first panel shows measurements in a random $X$ or $Z$ basis for each qubit. Optimization: A tensor network model is variationally optimized from $N$ samples using gradient descent, which outputs tensors $\hat{A}^{(N)}$ representing the state vector $|\psi(\hat{A}^{(N)})\rangle$. The middle panel illustrates an MPS model prior, as numerically investigated in this work. The loss function given by Eq. (5) includes the negative log-likelihood (NLL), typically used in maximum likelihood estimation (MLE). It is optionally regularized by physical observables estimated from the dataset (e.g., using classical shadow tomography represented by the pink dashed box). Evaluation: The algorithm outputs a tensor network or, in the current work, an MPS representation $\hat{A}^{(N)}$ as the learned state. We sometimes omit the number of samples and abbreviate the output tensor as $\hat{A}$. In the right panel, we evaluate the performance of our learning protocol by computing the fidelity $F=\lvert\mbox{$\langle\phi|\psi(\hat{A})\rangle$}\rvert$. From the reconstructed state, we can extract arbitrary physical observables, including the von Neumann entanglement entropy and highly nonlocal Pauli strings (the inset shows an example of the nonlocal $Z$ string operators considered in Fig. 4).
@@
To address the need for viable tomographic techniques for exotic two-dimensional systems [7], here, we demonstrate a heuristic state tomography method combining a tensor-network prior with maximum likelihood estimation (MLE) [26, 27, 28, 29] and randomized measurements [25, 30]. Previous works have considered learning one-dimensional quantum states from uniformly random single-qubit measurements [26] or two-dimensional quantum channels with random Paulli measurements [29]. In this work, we focus on two-dimensional real pure states and show that they are completely characterized by only randomized $XZ$ measurements, which randomly measure each qubit in either the $X$ or the $Z$ basis. The states considered here are not only potentially more complex than in previous studies but also easy to implement with high fidelity in experimental systems [31]. Moreover, we prove the tomographic completeness of randomized $XZ$ measurements on the space of pure and real quantum states, allowing the method to be extendable to nonstoquastic real Hamiltonians’ eigenstates [32]. Thus, our method provides a pragmatic pathway for characterizing two-dimensional states on near-term quantum devices, making use of only easy-to-implement measurements. We provide a detailed discussion of related works in Appendix A.
@@
In our numerical experiments, we use an MPS ansatz as our model prior. While MPS-based representations of two-dimensional states are inherently limited, our approach sets the stage for higher-dimensional tensor networks, such as projected entangled pair states (PEPS), for future investigations. In addition, our method provides a scalable approach that aligns with the existing capabilities of modern simulators, thereby providing a benchmark for more advanced tomographic techniques. To supplement our numerical results, we leverage the statistical properties of MLE to provide insights into the sample complexity of our approach. More specifically, we provide a pedagogical overview of standard techniques for analyzing MLE [33, 34, 35], and discuss where they fail in this context, due to gauge freedom in MPS representations of quantum states. We then sketch how these difficulties can be circumvented by considering MLE on a manifold [36], and apply these results to the variational manifold given by our MPS prior [37].
@@
The rest of this paper is structured as follows. In Sec. II, we present our protocol in conjunction with an in-depth discussion of our state-tomographic measurement schemes. Then, in Sec. III, we demonstrate the effectiveness of our protocol when applied to two topological states: a ground state of the surface code Hamiltonian and a $\mathbb{Z}_{2}$ quantum spin liquid state of Rydberg atoms arrayed on the ruby lattice. Our results are immediately relevant to current experiments since we only require feasible $XZ$ measurements and consider system sizes achievable with state-of-the-art simulators. Finally, we conclude by discussing the prospects and potential directions for improvement of our methodology in Sec. IV.
@@ II Variational tensor network tomography
In this section, we introduce our variational tensor network (TN) tomography protocol. To begin, we briefly review the framework of randomized measurements, wherein randomly chosen unitaries are applied to a target state before measuring it in the computational basis. Then, we utilize this data to variationally find the optimal tensor network approximation of the target state. We do so via maximum likelihood estimation, a standard technique for statistical inference; we also have the ability to regularize the MLE solution by estimations of physical quantities via classical shadows. This is done by adding a term to the MLE cost function. These steps are described below and represented by the schematic in Fig. 1.
@@
We follow the notation commonly used in randomized measurements [38, 25]. Let $|\phi\rangle$ be the unknown $n$-qubit pure state prepared by the quantum simulator, and let $\mathcal{U}$ be the unitary ensemble with a uniform distribution over a set $\{U\}$. We apply a unitary $U\sim\mathcal{U}$ before measuring in the computational basis. Performing this process repeatedly, we collect $N$ unitary-bit-string tuples as our dataset
@@
The unitary $U_{i}^{\alpha}$ rotates the measurement basis of the $i$-th qubit to the $\alpha=X,Z$ basis. Note that the most general global controls have been considered in earlier work [15], in which the same SU$(2)$ unitary is applied to each qubit. Here, we only focus on a discrete subset, of Pauli $X^{\otimes n}$ and $Z^{\otimes n}$ measurements, also considered in previous work [27]. For our second ensemble $\mathcal{U}_{XZ}$, the unitaries are uniformly sampled from
$\left\{\prod_{i}U_{i}^{\alpha},\,\alpha=X,Z\right\}.$ (4)
@@
Random-$XZ$ measurements $\mathcal{U}_{XZ}$ are tomographically complete on the space of real and pure states.
@@
To estimate the target state from the dataset $\{(U_{i},\,b_{i})\}_{i=1}^{N}$, we use a tensor network ansatz $|\psi({A})\rangle$, where a sequence of local tensors is denoted by ${A}=\left({A}^{[1]},\cdots,{A}^{[n]}\right)$. The advantage of this formulation is that one can efficiently compute the probability of a given measurement outcome in any basis by applying the corresponding unitary to the tensor network model as shown in the middle panel of Fig. 1. This allows us to evaluate the loss function, whose minimum leads to our estimate of the target state. The loss function is
@@
Figure 2: Numerical results for the surface code. (a) Schematic representation of the surface code of dimension $(L_{x},\,L_{y})$ with $Z$ stabilizers (red patches) and $X$ stabilizers (blue patches). The MPS “snakes” through the system along the path shown (black arrow). (b) Scaling of the infidelity with the number of samples $N$ for a $3\times 3$ surface code with global-$XZ$ (cross) or random-$XZ$ (circle) measurements. Shades of blue indicates the regularization strength $\beta$. (c) To investigate larger system sizes, we focus on the surface code in a strip geometry ($L_{y}=3$) with a perturbation ($h_{z}=0.1$). The three lines indicate the infidelity for different numbers of qubits, $n$, with random-$XZ$ measurements. Dark blue indicates a regularization $\beta=5$. The dashed line marks the fidelity threshold for $n=9$. (d) Sample complexity to reach a fixed local fidelity threshold $F_{\text{local}}=(0.99)^{n}$ extracted from (c). Colors and markers are the same as in (b).
@@
Matrix product states, which are one of the most extensively studied families of tensor networks, have enjoyed remarkable success in accurately representing many-body quantum states. Thanks to an effective variational algorithm known as the density-matrix renormalization group (DMRG) [40], MPS ansätze can be efficiently employed to describe the ground states of various many-body Hamiltonians. Strictly speaking, two-dimensional systems require the generalization of MPS, as the quasi-1D MPS description is limited to capturing short-range correlations and thus necessarily fails for sufficiently large and entangled systems. However, in practice, MPS ansätze are often used to represent ground states even in two dimensions, due to their relative advantage of efficient contractions. For the remainder of this paper, we utilize an MPS as our model prior, and we variationally find a state that is most consistent with the observed measurements, as shown in Fig. 1. The optimization of the MPS is carried out by taking gradients simultaneously with respect to all the tensor components; we provide the details in Appendix G.
@@ III.1 Learning the surface code
In our first demonstration, we consider learning the ground state of the perturbed surface code Hamiltonian [43], $H_{\text{sc}}=-\sum_{l}S_{Z,\,l}-\sum_{l}S_{X,\,l}-h_{z}\sum_{i}Z_{i}$, defined on the square lattice. Here, $S_{P,\,l}=\prod_{i\in\square_{l}}P_{i}$ denotes the $l$-th stabilizer, which is a product of Pauli $P$ operators ($P=X$ or $P=Z$) on a plaquette or boundary [Fig. 2(a)], and $h_{z}$ is a perturbative field. Such a perturbation drives the ground state away from the exactly solvable limit, but it can still be well approximated by a tensor network. Let $|\phi\rangle$ be the target MPS representation of the ground state along a snake path. Although our target state is possibly perturbed away from the exact stabilizer-state limit, the expectation values of the stabilizer operators should be smoothly connected to their fixed-point values ($\approx 1$) as a function of the perturbation strength. This knowledge informs our choice of the set of stabilizers as observables for regularization. We estimate these stabilizers from the finite dataset either via direct measurement in the global-$XZ$ scheme or from classical shadows for random-$XZ$ measurements.
Our results for the surface code are presented in Fig. 2. To begin, we consider the surface code without any external perturbation ($h_{z}=0$): for $n=9$ qubits arrayed on a square geometry, our protocol outputs the state vector $|\psi(\hat{A})\rangle$, and we evaluate the error of our learning protocol using the infidelity, $1-F=1-\lvert\mbox{$\langle\phi|\psi(\hat{A})\rangle$}\rvert$. Figure 2(b) shows that the infidelity improves with the number of samples $N$, for both the global-$XZ$ and random-$XZ$ measurements discussed in Sec. II. For the latter, we find that regularization over the stabilizers, by increasing $\beta$, further reduces the infidelity in the sample-limited regime ($N<1000$). We observe that both the global and random measurement schemes succeed in learning these states, and furthermore, the former outperforms the latter in terms of requiring fewer samples to achieve the same infidelity. However, while global-XZ measurements $\mathcal{U}_{\textnormal{Global}}$ are sufficient for certain highly structured states such as a surface code (or GHZ) state, they may not necessarily be suited for learning generic states, wherefore random-$XZ$ measurements $\mathcal{U}_{XZ}$ could be beneficial. Finally, we note that the infidelity scales polynomially with the number of samples $N$, similarly to observations in earlier work on one-dimensional states [26]. In particular, focusing on the exact surface code ground state (for $h_{z}=0$) trained without regularization ($\beta=0$), we find an excellent fit of $1-F\propto 1/N^{\alpha}$, as shown in Fig. 2(b), with $\alpha=1.16$ ($\alpha=1.05$) for random-$XZ$ (global-$XZ$) measurements. To substantiate these numerical observations regarding the sample complexity needed to achieve a certain fidelity, we rigorously establish in Theorem 2 that the infidelity can be probabilistically upper bounded such that
@@
Next, we add a perturbation to the Hamiltonian, $h_{z}=0.1$, and analyze the performance of our protocol for various system sizes with up to $n=45$ qubits. Given the limitations of an MPS ansatz, which is inherently quasi-one-dimensional, here, we only increase the length of the system $L_{x}$ along the $x$-axis. In Fig. 2(c), we illustrate how the infidelity scales for different system sizes (indicated by different lines) using random-$XZ$ measurements and find that the fidelity improves with $N$. Finally, we analyze how the sample complexity scales with the system size $n$. The requirement of achieving a constant fidelity is stringent as the overlap between two many-body states can vanish exponentially with $n$, a phenomenon known as the orthogonality catastrophe [44]. For a finite system size, however, the overlap still serves as a useful diagnostic indicator [45]. Therefore, we relax the requirement of a fixed global fidelity by considering instead the per-site local fidelity $F_{\text{local}}\equiv F^{1/n}$. Extrapolating from the data in (c), Fig. 2(d) plots the samples required to achieve a local fidelity threshold, e.g., as indicated by the dashed line in (c) for $n=9$ qubits. For the largest system size, $n=45$, we find that only a few thousand measurements are necessary to attain a fidelity of $F_{\text{local}}=0.99$.
Figure 3: Numerical results for learning quantum spin liquids. (a) The tuning parameters of the Rydberg Hamiltonian are the Rabi frequency $\Omega$, the detuning $\Delta$, and the blockade radius $R_{b}$. The gray arrow shows the mapping between a particular measurement bit-string on the ruby lattice and a dimer configuration on the kagome lattice: Light (dark) blue circles correspond to an atom being in the excited $|r\rangle$ (ground $|g\rangle$) state and maps to the presence (absence) of a dimer. A careful choice of the Rydberg blockade radius $R_{b}=2a$, where $a$ is the lattice spacing, ensures that the dimers do not overlap. Bottom panel: The two phases as a function of the detuning ratio $\Delta/\Omega$. (b) The ground state belongs to the trivial disordered phase at $\Delta/\Omega=0.5$. The infidelity decreases with an increasing number of samples $N$. For random-$XZ$ measurements with $N\leq 40000$, a regularization of $\beta=1$ (red circles) further improves the fidelity compared to the case without regularization $\beta=0$ (red squares). Global-$XZ$ measurements (gray crosses) perform worse than their random-$XZ$ counterparts (red). (c) Same as in (b) but for the spin-liquid state at $\Delta/\Omega=1.7$. Relative to the case with no regularization $\beta=0$ (light blue), regularization $\beta=1$ (dark blue) does not improve the fidelity but reduces the number of nonconvergent outliers (above the black dashed line). Global-$XZ$ (gray cross) measurements perform similarly to random-$XZ$ ones (blue).

### v4

@@ Abstract
Learning faithful representations of quantum states is crucial to fully characterizing the variety of many-body states created on quantum processors. While various tomographic methods such as classical shadow and MPS tomography have shown promise in characterizing a wide class of quantum states, they face unique limitations in detecting topologically ordered two-dimensional states. To address this problem, we implement and study a heuristic tomographic method that combines variational optimization on tensor networks with randomized measurement techniques. Using this approach, we demonstrate its ability to learn the ground state of the surface code Hamiltonian as well as an experimentally realizable quantum spin liquid state. In particular, we perform numerical experiments using MPS ansätze and systematically investigate the sample complexity required to achieve high fidelities for systems with sizes of up to $48$ qubits. In addition, we provide theoretical insights into the scaling of our learning algorithm by analyzing the statistical properties of maximum likelihood estimation. Notably, our method is sample-efficient and experimentally friendly, only requiring snapshots of the quantum state measured randomly in the $X$ or $Z$ bases. Using this subset of measurements, our approach can effectively learn any real pure states represented by tensor networks, and we rigorously prove that random-$XZ$ measurements are tomographically complete for such states.
@@ I Introduction
Tomographic and other recovery methods for extracting information from these experiments are crucial, in particular, when they provide detailed diagnostic information about the prepared quantum state [11]. Indeed, recent approaches using randomized measurements have shown promise for measuring specific properties, such as the Rényi entropy [8, 12, 13, 14]. However, extracting the von Neumann entanglement entropy across arbitrary subsystem sizes using these methods might require multiple copies, substantially more experimental control, or more samples than what is currently available. Quantum state tomography (QST) is designed to reconstruct a quantum state from informationally complete measurements and is useful for systems where the desired observables are not known a priori [15] or cannot be measured due to the constraints of the experimental setup. Although QST is generally not scalable [16, 17], it becomes significantly more effective when applied with priors for particular classes of states. This makes QST a valuable alternative for extracting many properties from such systems, particularly when traditional methods are limited. Established state tomographic methods like matrix product state (MPS) tomography [18, 19, 20] provide a natural description with robust guarantees for one-dimensional systems but fall short in accurately capturing two-dimensional quantum states, especially when they exhibit noninjective properties [21] and complex entanglement patterns [22]. While neural network quantum state tomography (NNQST) [23, 24] offers yet another option, its efficacy for topological states also remains unclear due to their intrinsic long-range entanglement and intricate phase structures across different components of the wavefunction [25].
@@
Figure 1: Schematic illustration of variational tensor network tomography. Data collection: A target state vector $|\phi\rangle$ is measured in multiple bases for a collection of $N$ samples defined in  Eq. (1). For instance, the first panel shows measurements in a random $X$ or $Z$ basis for each qubit. Optimization: A tensor-network model is variationally optimized from NN samples using gradient descent, which outputs tensors $\hat{A}^{(N)}$ representing the state vector $|\psi(\hat{A}^{(N)})\rangle$. The middle panel illustrates an MPS model prior, as numerically investigated in this work. The loss function given by Eq. (5) includes the negative log-likelihood (NLL), typically used in maximum likelihood estimation (MLE). It is optionally regularized by physical observables estimated from the dataset (e.g., using classical shadow tomography represented by the pink dashed box). Evaluation: The algorithm outputs a tensor network or, in the current work, an MPS representation $\hat{A}^{(N)}$ as the learned state. We sometimes omit the number of samples and abbreviate the output tensor as $\hat{A}$. In the right panel, we evaluate the performance of our learning protocol by computing the fidelity $F=\lvert\mbox{$\langle\phi|\psi(\hat{A})\rangle$}\rvert$. From the reconstructed state, we can extract arbitrary physical observables, including the von Neumann entanglement entropy and highly nonlocal Pauli strings (the inset shows an example of the nonlocal $Z$ string operators considered in Fig. 4).
@@
To address the need for viable tomographic techniques for exotic two-dimensional systems [7], here, we demonstrate a heuristic state tomography method combining a tensor-network prior with maximum likelihood estimation (MLE) [26, 27, 28, 29] and randomized measurements [25, 30]. Previous works have considered learning one-dimensional quantum states from uniformly random single-qubit measurements [26] or two-dimensional quantum channels with random Pauli measurements [29]. In this work, we focus on two-dimensional real pure states and show that they are completely characterized by only randomized $XZ$ measurements, which randomly measure each qubit in either the $X$ or the $Z$ basis. The states considered here are not only potentially more complex than in previous studies but also easy to implement with high fidelity in experimental systems [31]. Moreover, we prove the tomographic completeness of randomized $XZ$ measurements on the space of pure and real quantum states, allowing the method to be extendable to nonstoquastic real Hamiltonians’ eigenstates [32]. Thus, our method provides a pathway for characterizing two-dimensional states on near-term quantum devices, making use of only easy-to-implement measurements. For comparison, we also provide a detailed discussion of related works in Appendix A. While our main results assume perfect measurements, in Appendix H, we review the errors that are important when measuring states in Rydberg atom arrays and numerically investigate the effect of noise.
@@
In our numerical experiments, we use an MPS ansatz as our model prior. While MPS-based representations of two-dimensional states are inherently limited, our approach sets the stage for higher-dimensional tensor networks, such as projected entangled pair states (PEPS), for future investigations. In addition, our method provides a scalable approach that aligns with the existing capabilities of modern simulators, thereby providing a benchmark for more advanced tomographic techniques. To supplement our numerical results, we leverage the statistical properties of MLE to provide insights into the sample complexity of our approach. More specifically, we provide a pedagogical overview of standard techniques for analyzing MLE [33, 34, 35], and discuss where they fail in this context, due to the gauge freedom in MPS representations of quantum states. We then sketch how these difficulties can be circumvented by considering MLE on a manifold [36], and apply these results to the variational manifold given by our MPS prior [37].
@@
The rest of this work is structured as follows. In Sec. II, we present our protocol in conjunction with an in-depth discussion of our state-tomographic measurement schemes. Then, in Sec. III, we demonstrate the effectiveness of our protocol when applied to two topological states: a ground state of the surface code Hamiltonian and a $\mathbb{Z}_{2}$ quantum spin liquid state of Rydberg atoms arrayed on the ruby lattice. Our results are immediately relevant to current experiments since we only require feasible $XZ$ measurements and consider system sizes achievable with state-of-the-art simulators. Finally, we conclude by discussing the prospects and potential directions for improvement of our methodology in Sec. IV.
@@ II Variational tensor network tomography
In this section, we introduce our variational tensor network (TN) tomography protocol. To begin, we briefly review the framework of randomized measurements, wherein randomly chosen unitaries are applied to a target state before measuring it in the computational basis. Then, we utilize this data to variationally find the optimal tensor-network approximation of the target state. We do so via maximum likelihood estimation, a standard technique for statistical inference; we also have the ability to regularize the MLE solution with estimations of physical quantities via classical shadows. This is done by adding a term to the MLE cost function. These steps are described below and represented by the schematic in Fig. 1.
@@
We follow the notation commonly used in randomized measurements [38, 25]. Let $|\phi\rangle$ be the unknown $n$-qubit pure state prepared by the quantum simulator, and let $\mathcal{U}$ be the unitary ensemble with a uniform distribution over a set $\{U\}$. We apply a unitary $U\sim\mathcal{U}$ before measuring in the computational basis. Performing this process repeatedly, we collect $N$ unitary-bitstring tuples as our dataset
@@
The unitary $U_{i}^{a}$ rotates the measurement basis of the $i$-th qubit to the $a=X,Z$ basis. Note that the most general global controls have been considered in earlier work [15], in which the same SU(2)(2) unitary is applied to each qubit. Here, we only focus on a discrete subset, of Pauli $X^{\otimes n}$ and $Z^{\otimes n}$ measurements, also considered in previous work [27]. For our second ensemble $\mathcal{U}_{XZ}$, the unitaries are uniformly sampled from
$\left\{\prod_{i}U_{i}^{a},\,a=X,Z\right\}.$ (4)
@@
Random-XZXZ measurements $\mathcal{U}_{XZ}$ are tomographically complete on the space of real and pure states.
@@
To estimate the target state from the dataset $\{(U_{i},\,b_{i})\}_{i=1}^{N}$, we use a tensor-network ansatz $|\psi({A})\rangle$, where a sequence of local tensors is denoted by ${A}=\left({A}^{[1]},\cdots,{A}^{[n]}\right)$. The advantage of this formulation is that one can efficiently compute the probability of a given measurement outcome in any basis by applying the corresponding unitary to the tensor-network model as shown in the middle panel of Fig. 1. This allows us to evaluate the loss function, whose minimum leads to our estimate of the target state. The loss function is
@@
Figure 2: Numerical results for the surface code. (a) Schematic representation of the surface code of dimension $(L_{x},\,L_{y})$ with $Z$ stabilizers (red patches) and $X$ stabilizers (blue patches). The MPS “snakes” through the system along the path shown (black arrow). (b) Scaling of the infidelity with the number of samples $N$ for a $3\times 3$ surface code with global-XZXZ (cross) or random-XZXZ (circle) measurements. Shades of blue indicate the regularization strength $\beta$. (c) To investigate larger system sizes, we focus on the surface code in a strip geometry ($L_{y}=3$) with a perturbation ($h_{z}=0.1$). The three lines indicate the infidelity for different numbers of qubits, $n$, with random-$XZ$ measurements. Dark blue indicates a regularization $\beta=5$. The dashed line marks the fidelity threshold for $n=9$. (d) Sample complexity plotted on a loglog-loglog scale to reach a fixed local fidelity threshold $F_{\text{local}}=(0.99)^{n}$ extracted from (c). The number of samples required for achieving such a local fidelity is seen to scale at most polynomially with $n$. Colors and markers are the same as in (b).
@@
Matrix product states, which are one of the most extensively studied families of tensor networks, have enjoyed remarkable success in accurately representing many-body quantum states. Thanks to an effective variational algorithm known as the density-matrix renormalization group (DMRG) [40], MPS ansätze can be efficiently employed to describe the ground states of various many-body Hamiltonians. Strictly speaking, two-dimensional systems require the generalization of MPS, as the quasi-1D MPS description is limited to capturing short-range correlations and thus necessarily fails for sufficiently large and entangled systems. However, in practice, MPS ansätze are often used to represent ground states even in two dimensions, due to their relative advantage of efficient contractions. For the remainder of this work, we utilize an MPS as our model prior, and we variationally find a state that is most consistent with the observed measurements, as shown in Fig. 1. The optimization of the MPS is carried out by taking gradients simultaneously with respect to all the tensor components; we provide the details in Appendix G.
@@ III.1 Learning the surface code
In our first demonstration, we consider learning the ground state of the perturbed surface code Hamiltonian [43], $H_{\text{sc}}=-\sum_{l}S_{Z,\,l}-\sum_{l}S_{X,\,l}-h_{z}\sum_{i}Z_{i}$, defined on the square lattice. Here, $S_{P,\,l}=\prod_{i\in\square_{l}}P_{i}$ denotes the $l$-th stabilizer, which is a product of Pauli $P$ operators ($P=X$ or $P=Z$) on a plaquette or boundary [Fig. 2(a)], and $h_{z}$ is a perturbative field. Such a perturbation drives the ground state away from the exactly solvable limit, but it can still be well approximated by a tensor network. Let $|\phi\rangle$ be the target MPS representation of the ground state along a “snake” path. Although our target state is possibly perturbed away from the exact stabilizer-state limit, the expectation values of the stabilizer operators should be smoothly connected to their fixed-point values ($\approx 1$) as a function of the perturbation strength. This knowledge informs our choice of the set of stabilizers as observables for regularization. We estimate these stabilizers from the finite dataset either via direct measurement in the global-$XZ$ scheme or from classical shadows for random-$XZ$ measurements.
[…]
@@
Next, we add a perturbation to the Hamiltonian, $h_{z}=0.1$, and analyze the performance of our protocol for various system sizes with up to $n=45$ qubits. Given the limitations of an MPS ansatz, which is inherently quasi-one-dimensional, here, we only increase the length of the system $L_{x}$ along the $x$-axis. In Fig. 2(c), we illustrate how the infidelity scales for different system sizes (indicated by different lines) using random-$XZ$ measurements and find that the fidelity improves with $N$. Finally, we analyze how the sample complexity scales with the system size $n$. The requirement of achieving a constant fidelity is stringent as the overlap between two many-body states can vanish exponentially with $n$, a phenomenon known as the orthogonality catastrophe [44]. For a finite system size, however, the overlap still serves as a useful diagnostic indicator [45]. Therefore, we relax the requirement of a fixed global fidelity by considering instead the per-site local fidelity $F_{\text{local}}\equiv F^{1/n}$. Extrapolating from the data in (c), Fig. 2(d) plots the number of samples required to achieve a local fidelity threshold, e.g., as indicated by the dashed line in (c) for $n=9$ qubits. For the largest system size, $n=45$, we find that only a few thousand measurements are necessary to attain a fidelity of $F_{\text{local}}=0.99$.
Figure 3: Numerical results for learning quantum spin liquids. (a) The tuning parameters of the Rydberg Hamiltonian are the Rabi frequency $\Omega$, the detuning $\Delta$, and the blockade radius $R_{b}$. The gray arrow shows the mapping between a particular measurement bitstring on the ruby lattice and a dimer configuration on the kagome lattice: light (dark) blue circles correspond to an atom being in the excited $|r\rangle$ (ground $|g\rangle$) state and maps to the presence (absence) of a dimer. A careful choice of the Rydberg blockade radius $R_{b}=2a$, where $a$ is the lattice spacing, ensures that the dimers do not overlap. Bottom panel: The two phases as a function of the detuning ratio $\Delta/\Omega$. (b) The ground state belongs to the trivial disordered phase at $\Delta/\Omega=0.5$. The infidelity decreases with an increasing number of samples $N$. For random-XZXZ measurements with $N\leq 40000$, a regularization of $\beta=1$ (red circles) further improves the fidelity compared to the case without regularization $\beta=0$ (red squares). Global-XZXZ measurements (gray crosses) perform worse than their random-XZXZ counterparts (red). (c) Same as in (b) but for the spin-liquid state at $\Delta/\Omega=1.7$. Relative to the case with no regularization $\beta=0$ (light blue), regularization $\beta=1$ (dark blue) does not improve the fidelity but reduces the number of nonconvergent outliers (above the black dashed line). Global-XZXZ (gray cross) measurements perform similarly to random-XZXZ ones (blue).